# Enhancing universal machine learning potentials with polarizable long-range interactions

### v1

@@ Results
To detect subtle lattice distortions and variations in free energy that distinguish different phases, a 10×10×10 supercell of $\rm BaTiO_{3}$ was simulated. This allowed for a detailed analysis of the structural changes that occurred between different phases, providing valuable insights into the properties of materials. Fig. 3a and b illustrate the simulation results, which clearly identify four distinct phases and the three first-order phase transitions within the simulated temperature range. Below 145 K, the overall average polarization of the supercell aligns with the [111] direction, indicative of the rhombohedral phase. At 145 K, a phase transition to orthorhombic is suggested as the y-component of polarization approaches zero. With further increase to 205 K, the polarization predominantly orients along the x-direction, indicating the tetragonal phase. This phase persists until 285 K, where the cubic paraelectric phase is observed. The obtained phase transition temperatures align closely with those predicted by various models and experiments, as shown in Table 1. Notably, utilizing first principles methods or fitting potential energy surfaces to first principles data often underestimates these temperatures, which is attributed to the approximated exchange-correlation functional45 employed in DFT. Although the predicted phase transition temperatures are lower than the experimentally observed values, the model effectively captures the sequence of phase transitions as measured in experiments, demonstrating the efficacy of the pretrained universal model in modeling highly polarized phase transitions in ferroelectric materials. In addition, we investigated the response to external electric fields, focusing specifically on the hysteresis behavior of $\rm BaTiO_{3}$. By applying varying electric fields, we can chart the polarization-electric field hysteresis loop (P-E loop). Experimental studies have shown that the critical thickness of $\rm BaTiO_{3}$ for ferroelectricity is about 4.8 nm48. To clearly illustrate the ferroelectric phenomenon, we employed a supercell of 20×10×10 tetragonal $\rm BaTiO_{3}$ ($\sim 8.2$ nm along a-direction) to simulate the P-E loop at 235 K. The electric field was applied along the x-direction, and the resulting changes in the x-component local polarization of the unit cell are depicted in Fig. 3c. The hysteresis loop obtained from our simulation reveals that the coercive electric field $E_{c}$ is about 300 kV/cm. This finding is consistent with experimental measurement of thin film $\rm BaTiO_{3}$49, 50 and effective Hamiltonian calculations51, both of which report values ranging from 200-500 kV/cm. Our universal model excels in temperature-dependence phase transitions and responses to external electric fields in ferroelectric materials. This proficiency holds significant promise for exploring ferroelectric characteristics, which are crucial for the development of non-volatile memory devices, transducers, microsensors, and capacitors.
@@
This work 145 205 285
First Principles52 150±10 195±5 265±5
Second Principles53 140 180 224
Effective Hamiltonian44 119 158 257
ReaxFF54 N.A. N.A. 240
MLIPs42 18.6 91.4 182.4
Experiments55 183 278 403
Interface formation of lithium thiophosphates
Figure 4: The formation of the lithium thiophosphates and lithium metal SEI layer at 300 K and 1 atm. a, The initial structure of $\beta-\rm{Li_{3}PS_{4}}$ (010)/Li (001) interface. The elements are color-coded: Li in green, S in purple, and P in pink. A segment of the primary reaction zone, ranging from 8 to 26 nanometers is displayed. b, Snapshots at different times during the MD simulations. The partial charges on lithium ions are represented with color coding to enhance the visibility of structural transformations during the formation of SEI. The SEI layer, approximately 8.5 nm in thickness, forms after 4 ns MD simulations, comprising an amorphous $\beta-\rm{Li_{3}PS_{4}}$/Li2S interface ($\sim 2$ nm), a crystalline Li2S layer ($\sim 4.5$ nm), and an amorphous Li2S (Li3P)/Li interface layer ($\sim 2$ nm) in sequence. c, The distributions of the Li partial atomic charges along the x-direction at 10, 100, 2000, and 4000 ps. d, The radial distribution function plots of Li-S and Li-P within the electrolyte and SEI layer.
Figure 5: The decomposition dynamics of Li3PS4 electrolyte and component analyses of SEI. The snapshots capture the dynamic decomposition behavior of PS4, showcasing: a, the sulfur, and b, the phosphorus component. To elucidate the formation mechanism of the SEI layer , atomic component analyses of the c, crystalline, and d, amorphous Li2S regions are compared among 4-ns (upper), 2-ns (middle), and the initial state (lower). The elements are color-coded: Li in green, S in purple, and P in pink. The x-component root mean squared displacements (RMSDx) of Li, P, and S between the initial and the 2-ns states and between 4-ns and 2-ns states are also shown.
@@
The final application studied in this work involves all-solid-state batteries, which represent a breakthrough in the evolution of next-generation energy solutions, owing to their superior energy density and inherent safety features56. Lithium thiophosphate, known for its exceptional ionic conductivity ($\sim 10^{-3}$ S/cm), is deemed the most promising candidate for solid electrolytes and has been widely studied through experiments57, 58, 59 and computer modelings60, 61, 62. Notably, the nanoporous $\beta-\rm{Li_{3}PS_{4}}$ has been validated to exhibit outstanding cycling stability63, presumably attributed to the formation of Li2S and Li3P solid-electrolyte interphases (SEI) during the initial battery cycles, which serves to passivate further degradation of the electrolyte61, 64, 65, 66. To investigate the formation of SEI, we utilized the pretrained universal model to simulate the interfacial MD of the solid electrolyte $\beta-\rm{Li_{3}PS_{4}}$ in contact with the lithium metal anode, encompassing a system of 13,760 atoms.
@@
The initial $\beta-\rm{Li_{3}PS_{4}}$/Li interfacial structure, measuring 33.15 nm × 3.1 nm × 2.6 nm, was utilized in the MD simulations, as depicted in Fig. 4a. The highly reactive Li metal began to interact with PS4 tetrahedron at the interface, triggering the formation of the SEI layer as illustrated in Fig. 4b. With the growth of the SEI layer, electrons were transferred from the Li anode to the $\rm{Li_{3}PS_{4}}$ electrolyte, which causes a gradual increase of the partial charges of Li as they migrate from the anode, across the SEI, and into the electrolyte. This results in a transition from metal lithium to ions. The dynamic behavior of partial charges at the anode-electrolyte interface is a critical aspect showcased in Fig. 4c. Initially, at 10 ps, the partial charges of lithium changed almost linearly from the anode to the electrolyte. Over time, this distribution evolved and showed a distinct plateau at 100 ps. This phenomenon could be attributed to the ordered structuring of the SEI, signifying the nucleation of crystalline structures. Ultimately, a stable lithium partial charges plateau was formed between the anode and the electrolyte, spanning a range of 15 to 19.5 nm. This is further supported by the visualized structure depicted in Fig. 4b, where the plateau within the SEI aligns with the Li2S crystal region. Chen et al.60 employed AIMD simulations to investigate the radial distribution functions (RDF) of Li-P and Li-S bonds, which concluded that the decompositions of electrolytes in lithium thiophosphates are primarily due to the decomposition of PS4 tetrahedron by the active lithium metal, leading to the formation of new Li-P and Li-S bonds. Our RDF analysis confirmed the formation of Li-S and Li-P bonds within the SEI and electrolyte as depicted in Fig. 4d. In agreement with the AIMD simulations60, we observed the emergence of a new Li-P peak within the SEI at approximately 2.5 Å. Furthermore, within the SEI, the Li-S bond displayed RDF characteristics consistent with crystalline Li2S67. For the Li-P bonds, the peak shape closely resembles that of amorphous Li3P AIMD simulations68. As illustrated in Fig. 5a and b, the ultimate decomposition products of P form only short-range ordered structures, in contrast to S, which forms long-range ordered structures. This observation agrees with the measurements where Li2S crystals were detected but Li3P crystals were absent in cryogenic transmission electron microscopy (cryo-TEM) experiments, as noted in reference58. To enhance the understanding of the formation mechanism, the atomic compositions of the SEI in both crystalline and amorphous Li2S regions were examined. As depicted in Fig. 5c and d, within the first 2 ns, the emergence of amorphous Li2S and Li3P regions was attributed to the swift diffusion of P and S atoms. Conversely, the formation of crystalline Li2S regions was predominantly influenced by the swift diffusion of Li atoms and slower diffusion of P and S atoms. As the system further evolves to 4 ns, the diffusions slow down, which also confirmed that the formation of crystalline Li2S region hindered further atomic diffusion and thus slowed the growth of the SEI. As illustrated in Fig. 4b, during the initial phase of decomposition (10-100 ps), the swift diffusion of P, S, and Li led to the formation of an interface of about 3 nm thick. However, in a later stage between 100 and 2000 ps, the formation of Li2S nuclei impeded the diffusion of P and S atoms, resulting in a period of sluggish SEI growth. Consequently, the interface experienced only a slight expansion in this period, increasing by only 1 nm. Furthermore, in the following 2 ns, there was essentially no SEI growth. The final interface structure ($\sim 8.5$ nm) could be characterized by a 4.5 nm crystalline Li2S region sandwiched between two 2-nm transitional layers. The cessation of growth at this interface suggests that the crystalline Li2S and amorphous Li3P within the crystal region contribute to the stabilization of both the lithium metal anode and the electrolyte, which agrees with the theoretical predictions made by DFT calculations64. Additionally, we studied the interface reactions between $\rm Li_{6}PS_{5}Cl$ and lithium metal, as detailed in Supplementary Information Part 4. As shown in Fig. S1, at a pressure of 1 atm and temperature of 300 K, $\rm Li_{6}PS_{5}Cl$ formed an interface of about 5-nm thick after 2 ns. This indicates that the $\rm Li_{6}PS_{5}Cl$ electrolyte exhibits a higher stability against lithium metal at room temperature compared to $\rm Li_{3}PS_{4}$. Upon elevating the temperature to 500 K, polycrystalline Li2S was detected, in agreement with cryo-TEM observations58.
The above molecular dynamics simulations highlight the superiority of our pretrained model in describing the interfacial reactions, which will significantly advance the development of future all-solid-state battery applications. This approach allows for a detailed examination of the interactions and behaviors at the atomic level. Through these simulations, valuable insights can be gained into the dynamic interactions that govern the stability and performance of all-solid-state batteries.
@@ Discussion
In this study, we introduced a novel framework that integrates short-range equivariant machine learning interatomic potentials with long-range polarizable interactions. This framework enhances the precision of long-range interactions by dynamically minimizing potential energy with respect to charge fluctuations, effectively incorporating charge transfer and polarization effects. By utilizing electrostatic interaction energy directly instead of relying on partial charges, the proposed framework facilitates the development of a universal model with broad applicability. The robustness of the model is further demonstrated through benchmarking of mechanical properties, specifically the prediction of bulk modulus across a large dataset of materials. The high $R^{2}$ values achieved indicate reliability of the model in predicting material properties that were not explicitly included in the training process. This capability is essential for high-throughput screening, potentially accelerating the discovery of materials with exceptional mechanical properties. In addition, the model has been successfully applied to study dynamic properties, such as lithium-ion diffusion in solid-state electrolytes and phase transitions in ferroelectric materials. Its ability to replicate AIMD simulations, while significantly reducing computational costs, suggests that it could be a valuable tool for investigating kinetic properties across a wide range of materials. Moreover, the simulations of temperature-dependence phase transition of ferroelectric material BaTiO3, as well as its P-E loop confirmed the capability of the model in addressing polarizable systems. Regarding the interfacial reactions between lithium thiophosphate electrolytes and lithium metal anodes in solid-state batteries, the model successfully elucidated the mechanism of interface formation and its impact on battery performance, highlighting its potential for realistic-scale simulations.
In conclusion, the pretrained universal model presented in this work not only delivers commendable performance across various domains but also establishes a solid foundation for future high-throughput material screening and a deeper understanding of material behaviors. This model unlocks invaluable insights that have the potential to propel the development of a new generation of high-performance materials and transformative energy solutions. The implications of this work are far-reaching and significantly enhance our technological capabilities.
@@ Methods
Dataset
To train a universal equivariance neural network potential, we used the MPtrj Dataset16 sourced from Materials Projects35 as the pretraining dataset. All configurations were calculated using DFT with the PBE69/PBE+U70 exchange-correlation functional and pseudopotential basis. The PBE/PBE+U mixing compatibility correction71 was applied to ensure energy consistency. After removing unreasonable configurations, the dataset comprised a total of 1,557,645 material configurations, each with corresponding energies, forces, and stresses.
@@
The higher-order energy terms and the electrostatic Coulomb interaction energies due to charge fluctuations should be deducted from the total potential energy. Appropriately setting the PQEq parameters is crucial for training a universal framework. In particular, the Li parameters were refitted to reproduce the electrostatic interaction energies observed in QM. Detailed information on the fitting process and the PQEq parameters used in this work can be found in the Supplementary Information Part 2.
For equivariant neural network training, the NequIP12 model was utilized, incorporating six equivariant layers with node features set at 256 × 0e + 256 × 1e. Spherical harmonic basis functions were used to represent interatomic directions, denoted as 1 × 0e + 1 × 1o. Interatomic distances were described using eight Bessel basis functions, and a cut-off value of 5.0 Å was chosen for building the short-range neighbor list. The dataset is divided into training set, validation set, and test set in a ratio of 6:1:1.
@@
where energies $\bm{E}$, forces $\bm{F}$ and stresses $\bm{S}$ weights were set as $\omega_{E}=1$, $\omega_{F}=10$, and $\omega_{S}=10$ with quantity units of eV, eV/Å, and eV/Å3. The batch size was selected as 128, and the Adam optimizer was employed with an initial learning rate of 10-3. A total of 71,452,040 trainable parameters were initialized with the random seed 3407. The implementation of the entire neural networks and long-range interactions was based on JAX 0.4.2072. The training was conducted on a single NVIDIA H100-PCIe GPU, utilizing CUDA version 12.3 and Driver version 545.23.06.
Dynamics
@@
The Atomic Simulation Environment (ASE)73 was utilized as the interface for geometry relaxations and MD simulations. Structure relaxations were conducted using the limited-memory BFGS74 method. Both atomic coordinates and cell vectors were optimized concurrently until the forces fell below the convergence threshold of 0.05 eV/Å.
@@
The c-LLZO with a space group of $\rm Ia\bar{3}d$ originated from experimental structures detailed in reference39. The structure was expanded to a 2×2×2 supercell. Subsequently, 180 lithium atoms were randomly placed on the 24(d) sites, and an additional 268 lithium atoms on the 96(h) sites. After relaxation, the structure served as the starting point for MD simulations. The isothermal isobaric (NpT) ensemble simulations were performed with the temperature control via Nosé-Hoover thermostat75, 76 and pressure maintained at 1 atm using Parrinello-Rahman barostat77, 78 to determine the lattice constants at various temperatures. Then, the canonical (NVT) ensemble simulations with the Nosé-Hoover thermostat were utilized to ascertain the lithium diffusion properties within c-LLZO. A time step of 2 fs was used for all NpT and NVT simulations. In the case of NVT simulations, following a 100 ps equilibration period, 2-ns trajectories’ mean squared displacements of Li ions were employed to calculate the self-diffusion coefficients at various target temperatures. Additionally, we conducted an uncertainty analysis of the diffusion coefficients, applying the empirical error estimation method as detailed in reference34.
[…]
@@
We utilized a 13760-atom structure consisting of a ca. 13.6-nm thick $\rm Im\bar{3}m$ lithium metal anode with [100] facet, paired with a ca. 19.1 nm thick $\rm Pnma$ $\beta-\rm{Li_{3}PS_{4}}$ electrolyte with [010] facet to illustrate the SEI formation across realistic spatial and temporal scales. The system was structured with periodic lateral dimensions measuring 3.1 nm by 2.6 nm. The unit cell of anode and electrolyte structures originated from the Materials Project35. The entire cell was subject to periodic boundary conditions, with a 1 nm atomic layer fixed on each side along the x-direction to ensure a singular interface reaction. NpT MD simulations were conducted at a standard temperature of 300 K and a pressure of 1 atm by Nosé-Hoover thermostat and Parrinello-Rahman barostat with a timestep of 2 fs. Initial velocities for the particles were assigned randomly following the Maxwell-Boltzmann distribution, and the simulations were carried out for a duration of 4 ns.
Data availability
@@
Source data will be provided with this paper. The dataset is accessible through the Materials Project35, 16. Detailed PQEq parameters are provided in the Supplementary Information.
Code availability
The NequIP12 code originated from the open-source project jax-md79. The JAX implementation of PQEq and DFT-D3 will be made available at https://github.com/reaxnet/jax-nb.

### v2

@@ Results
To verify the ability of our model to handle polarized systems, we performed MD simulations on a typical perovskite ferroelectric material $\rm BaTiO_{3}$. As temperature increases, $\rm BaTiO_{3}$ undergoes a non-trivial phase transition process44. Below 183 K, it adopts a rhombohedral structure with local polarization along the [111] direction. As the temperature rises, it transitions to an orthorhombic phase with polarization along the [110] direction, then to a tetragonal phase with polarization along [100] direction at 278 K. Finally, at 403 K, it converts to a cubic paraelectric phase. Accurately simulating phase changes in ferroelectric substances requires precise potential energy functions that can respond to the small atomic shifts and structural changes, as well as account for the free energy surfaces under the conditions of finite-temperature thermodynamics45. The rhombohedral-orthorhombic-tetragonal-cubic phase sequence of $\rm BaTiO_{3}$ has been extensively studied and reproduced using effective Hamiltonians46, 47, CMD48, 49, and specialized MLIPs models445, 50. To detect subtle lattice distortions and variations in free energy that distinguish different phases, a 10×10×10 supercell of $\rm BaTiO_{3}$ was simulated. This allowed for a detailed analysis of the structural changes that occurred between different phases, providing valuable insights into the properties of materials. The use of larger simulation cells (10×10×10 in our work versus 4×4×4 in Gigli et al. 45) may provide some advantages for phase transition studies. Larger supercells can effectively reduce the impact of temperature fluctuations, enabling better temperature control for sampling the free energy landscape, thereby yielding results with greater statistical significance. Fig. 4 a and b illustrate the simulation results, which clearly identify four distinct phases and the three first-order phase transitions within the simulated temperature range. Below 145 K, the overall average polarization of the supercell aligns with the [111] direction, indicative of the rhombohedral phase. At 145 K, a phase transition to orthorhombic is suggested as the y-component of polarization approaches zero. With further increase to 205 K, the polarization predominantly orients along the x-direction, indicating the tetragonal phase. This phase persists until 285 K, where the cubic paraelectric phase is observed. The obtained phase transition temperatures align closely with those predicted by various models and experiments, as shown in Table 2. Notably, utilizing first principles methods or fitting potential energy surfaces to first principles data often underestimates these temperatures, which is attributed to the approximated exchange-correlation functional48 employed in DFT. Although the predicted phase transition temperatures are lower than the experimentally observed values, the model effectively captures the sequence of phase transitions as measured in experiments, demonstrating the efficacy of the foundation model in modeling highly polarized phase transitions in ferroelectric materials.
Figure 4: The temperature dependence of a, lattice constants, and b, local polarizations of unit cells exhibit notable changes during the phase transitions observed from MD simulations on 10×10×10 supercell of BaTiO3.
@@
This work (10×10×10) 145 205 285
Effective Hamiltonian (12×12×12) 51 150±10 195±5 265±5
Effective Hamiltonian (16×16×16) 47 119 158 257
Second Principles (16×16×16) 52 140 180 224
ReaxFF (6×6×6) 53 N.A. N.A. 240
MLIPs (4×4×4) 45 18.6 91.4 182.4
Experiments54 183 278 403
@@
The final application studied in this work involves all-solid-state batteries, which represent a breakthrough in the evolution of next-generation energy solutions, owing to their superior energy density and inherent safety features55. Lithium thiophosphate, known for its exceptional ionic conductivity ($\sim 10^{-3}$ S/cm), is deemed the most promising candidate for solid electrolytes and has been widely studied through experiments56, 57, 58 and computer modelings59, 60, 61. Notably, the nanoporous $\beta-\rm{Li_{3}PS_{4}}$ has been validated to exhibit outstanding cycling stability62, presumably attributed to the formation of Li2S and Li3P solid-electrolyte interphases (SEI) during the initial battery cycles, which serves to passivate further degradation of the electrolyte60, 63, 64, 65. To investigate the formation of SEI, we utilized the foundation model to simulate the interfacial MD of the solid electrolyte $\beta-\rm{Li_{3}PS_{4}}$ in contact with the lithium metal anode, encompassing a system of 13,760 atoms.
Figure 5: a, The initial structure of $\beta$-Li3PS4 [010]/Li [001] interface. The elements are colour-coded: Li in green, S in purple, and P in pink. A segment of the primary reaction zone, ranging from 8 to 26 nanometres is displayed. b, Snapshots at different times during the MD simulations. The partial charges on lithium ions are represented with colour coding to enhance the visibility of structural transformations during the formation of SEI. The SEI layer, approximately 8.5 nm in thickness, forms after 4 ns MD simulations, comprising an amorphous $\beta$-Li3PS4/Li2S interface ( 2 nm), a crystalline Li2S layer ( 4.5 nm), and an amorphous Li2S (Li3P)/Li interface layer ( 2 nm) in sequence. c, The distributions of the Li partial atomic charges along the x-direction at initial, 20, 100, 2000, and 4000 ps state. d, The radial distribution function plots of Li-S and Li-P within the electrolyte and SEI layer.
Figure 6: The snapshots capture the dynamic decomposition behavior of PS4, showcasing: a, the sulfur, and b, the phosphorus component. To elucidate the formation mechanism of the SEI layer , atomic component analyses of the c, crystalline, and d, amorphous Li2S regions are compared among 4-ns (upper), 2-ns (middle), and the initial state (lower). The elements are color-coded: Li in green, S in purple, and P in pink. The x-component root mean squared displacements (RMSDx) of Li, P, and S between the initial and the 2-ns states and between 4-ns and 2-ns states are also shown.
@@
The initial $\beta-\rm{Li_{3}PS_{4}}$/Li interfacial structure, measuring 33.15 nm × 3.1 nm × 2.6 nm, was utilized in the MD simulations, as depicted in Fig. 5 a. The highly reactive Li metal began to interact with PS4 tetrahedron at the interface, triggering the formation of the SEI layer as illustrated in Fig. 5 b. With the growth of the SEI layer, electrons were transferred from the Li anode to the $\rm{Li_{3}PS_{4}}$ electrolyte, which causes a gradual increase of the partial charges of Li as they migrate from the anode, across the SEI, and into the electrolyte. This results in a transition from metal lithium to ions. The dynamic behavior of partial charges at the anode-electrolyte interface is a critical aspect showcased in Fig. 5 c. Initially, at 10 ps, the partial charges of lithium changed almost linearly from the anode to the electrolyte. Over time, this distribution evolved and showed a distinct plateau at 100 ps. This phenomenon could be attributed to the ordered structuring of the SEI, signifying the nucleation of crystalline structures. Ultimately, a stable lithium partial charges plateau was formed between the anode and the electrolyte, spanning a range of 15 to 21.5 nm. This is further supported by the visualized structure depicted in Fig. 5 b, where the plateau within the SEI aligns with the Li2S crystal region. Chen et al.59 employed AIMD simulations to investigate the radial distribution functions (RDF) of Li-P and Li-S bonds, which concluded that the decompositions of electrolytes in lithium thiophosphates are primarily due to the decomposition of PS4 tetrahedron by the active lithium metal, leading to the formation of new Li-P and Li-S bonds. Our RDF analysis confirmed the formation of Li-S and Li-P bonds within the SEI and electrolyte as depicted in Fig. 5 d. In agreement with the AIMD simulations59, we observed the emergence of a new Li-P peak within the SEI at approximately 2.5 Å. Furthermore, within the SEI, the Li-S bond displayed RDF characteristics consistent with crystalline Li2S66. For the Li-P bonds, the peak shape closely resembles that of amorphous Li3P AIMD simulations67. As illustrated in Fig. 6 a and b, the ultimate decomposition products of P form only short-range ordered structures, in contrast to S, which forms long-range ordered structures. This observation agrees with the measurements where Li2S crystals were detected but Li3P crystals were absent in cryogenic transmission electron microscopy (cryo-TEM) experiments, as noted in reference57. To enhance the understanding of the formation mechanism, the atomic compositions of the SEI in both crystalline and amorphous Li2S regions were examined.
As depicted in Fig. 6 c and d, within the first 2 ns, the emergence of amorphous Li2S and Li3P regions was attributed to the swift diffusion of P and S atoms. Conversely, the formation of crystalline Li2S regions was predominantly influenced by the swift diffusion of Li atoms and slower diffusion of P and S atoms. As the system further evolves to 4 ns, the diffusions slow down, which also confirmed that the formation of crystalline Li2S region hindered further atomic diffusion and thus slowed the growth of the SEI. As illustrated in Fig. 5 b, during the initial phase of decomposition (10-100 ps), the swift diffusion of P, S, and Li led to the formation of an interface of about 3 nm thick. However, in a later stage between 100 and 2000 ps, the formation of Li2S nuclei impeded the diffusion of P and S atoms, resulting in a period of sluggish SEI growth. Consequently, the interface experienced only a slight expansion in this period, increasing by only 1 nm. Furthermore, in the following 2 ns, there was essentially no SEI growth. The final interface structure ($\sim 8.5$ nm) could be characterized by a 4.5 nm crystalline Li2S region sandwiched between two 2-nm transitional layers. The cessation of growth at this interface suggests that the crystalline Li2S and amorphous Li3P within the crystal region contribute to the stabilization of both the lithium metal anode and the electrolyte, which agrees with the theoretical predictions made by DFT calculations63.
Additionally, we studied the interface reactions between $\rm Li_{6}PS_{5}Cl$ and lithium metal, as detailed in Supplementary Information. As shown in Fig. S12, at a pressure of 1 atm and temperature of 300 K, $\rm Li_{6}PS_{5}Cl$ formed an interface of about 5-nm thick after 2 ns. This indicates that the $\rm Li_{6}PS_{5}Cl$ electrolyte exhibits a higher stability against lithium metal at room temperature compared to $\rm Li_{3}PS_{4}$. Upon elevating the temperature to 500 K, we detected Li2S at the grain boundaries, which aligns with the cryo-TEM observations 57.
Finetuning for enhanced accuracy
[…]
@@ Discussion
In this study, we introduced a novel framework that integrates equivariant machine learning interatomic potentials with explicit long-range polarizable interactions. QEq-based models inherently struggle with atomic polarization because they do not distinguish between core and shell components. Due to their inability to handle atomic-level responses to external fields, they may exhibit non-physical behaviours in some cases as described in the reference 35. Our framework incorporates polarizable long-range interactions and accurately captures atomic polarization in the presence of external electric fields. Since our framework decomposes the total energy into electrostatic energy and remaining energies that only depend on atomic types and positions, any future machine learning techniques and/or high-order electrostatic energy expansion schemes can be readily integrated into it, empowering possible accuracy enhancement.
Our foundation model successfully reproduced DFT potential energy surfaces for Na-Cl/Na-Na dimers across extended interatomic distances, outperforming conventional message-passing MLIP beyond the cut-off regions. The model demonstrates exceptional versatility across multiple materials science challenges, including accurate prediction of thermal stability in complex layered materials, high-fidelity capture of molecular responses to external electric fields, and reliable prediction of bulk modulus. The model has been also successfully applied to study dynamic properties and reactive molecular dynamics, such as lithium-ion diffusion in solid-state electrolytes and temperature-dependent phase transitions in ferroelectric materials like BaTiO3, while also elucidating interface formation mechanisms between lithium thiophosphate electrolytes and lithium metal anodes in solid-state batteries.
Beyond these specific applications, we evaluated the model’s transferability, a critical attribute for a foundation model. The model accurately reproduces the potential energy surfaces for OH-OH systems, correctly distinguishing between the neutral state and the long-range Coulombic repulsion of the ionic state. The model achieved a remarkable mean absolute error of just 4.57 a.u. for molecular polarizability, highlighting its robust and generalizable learned representation of polarization physics. Finally, we demonstrated our foundation model’s capacity to achieve ab initio accuracy for specific systems through efficient finetuning.
We acknowledge that our model aims to achieve a balance between computational efficiency and accuracy. Several advanced machine learning approaches for long-range interactions may achieve higher accuracy for specific, complex systems, including higher-order charge expansion methods and simultaneous optimization of QEq parameters 68, 69 in each step. However, more accurate methods incur higher computational complexity. The method that dynamically solve charge equilibration parameters at each step require additional cost for MD simulations. It is also observed that the charge equilibration parameters do not change significantly during MD simulations for that method 69, suggesting that expanding polarization charges to higher orders while using fixed, a priori parameters may achieve higher accuracy with tolerable computational overheads. Additionally, several approaches that do not require charge equilibration process provide valuable insights. Latent Ewald Summation 20 captures long-range interactions by learning hidden variables from local descriptors and applying Ewald summation to these variable. LSR-MP 70 employs a fragmentation-based approach with hierarchical message passing between atoms and fragments to learn the long-range interactions. NeuralP3M 71 introduces mesh points alongside atoms to discretize space and capture long-range effects. SO3krates 72 utilizes spherical harmonic coordinates and equivariant attention for interactions at arbitrary length scales. In contrast to these approaches, our foundation model explicitly simulates charge redistribution through a physics-driven framework that maintains a transparent connection to the underlying physical principles of charge polarization and electrostatic interactions. This design choice enhances interpretability while preserving computational efficiency. A promising direction for future work lies in exploring strategic integrations of our framework with these complementary methodologies to develop even more comprehensive and powerful foundation models.
@@ Methods
To train the foundation equivariance neural network potential, we used the MPtrj Dataset16 sourced from Materials Projects36 as the training dataset. All configurations were calculated using DFT with the PBE73/PBE+U74 exchange-correlation functional and pseudopotential basis. The PBE/PBE+U mixing compatibility correction75 was applied to ensure energy consistency. Due to transferability issues with the Yb element 36, all data containing Yb in the MPtrj dataset were excluded. The higher-order energy terms and the electrostatic Coulomb interaction energies due to charge fluctuations should be deducted from the total potential energy. Appropriately setting the PQEq parameters is crucial for training a universal framework. In particular, the Li parameters were refitted to reproduce the electrostatic interaction energies observed in QM. Detailed information on the fitting process and the PQEq parameters used in this work can be found in the Supplementary Information.
For equivariant neural network training, the NequIP12 model was utilized, incorporating six equivariant layers with node features set at 256 × 0e + 256 × 1e. Spherical harmonic basis functions were used to represent interatomic directions, denoted as 1 × 0e + 1 × 1o. Interatomic distances were described using eight Bessel basis functions, and a cut-off value of 5.0 Å was chosen for building the short-range neighbor list. The dataset is divided into training set, validation set, and test set in a ratio of 8:1:1.
@@
where energies $\bm{E}$, forces $\bm{F}$ and stresses $\bm{S}$ weights were set as $\omega_{E}=1$, $\omega_{F}=10$, and $\omega_{S}=10$ with quantity units of eV, eV/Å, and eV/Å3. The batch size was selected as 128, and the Adam optimizer was employed with an initial learning rate of 10-3. Trainable parameters were initialized with the random seed 3407. The implementation of the entire neural networks and long-range interactions was based on JAX 0.4.2076. The training was conducted on a single NVIDIA H100-PCIe GPU, utilizing CUDA version 12.3 and Driver version 545.23.06.
Molecular dynamics simulations
@@
The Atomic Simulation Environment (ASE)77 was utilized as the interface for geometry relaxations and MD simulations. Structure relaxations were conducted using the limited-memory BFGS78 method. Both atomic coordinates and cell vectors were optimized concurrently until the forces fell below the convergence threshold of 0.05 eV/Å.
@@
The c-LLZO with a space group of $\rm Ia\bar{3}d$ originated from experimental structures detailed in reference42. The structure was expanded to a 2×2×2 supercell. Subsequently, 180 lithium atoms were randomly placed on the 24(d) sites, and an additional 268 lithium atoms on the 96(h) sites. After relaxation, the structure served as the starting point for MD simulations. The isothermal isobaric (NpT) ensemble simulations were performed with the temperature control via Nosé-Hoover thermostat79, 80 and pressure maintained at 1 atm using Parrinello-Rahman barostat81, 82 to determine the lattice constants at various temperatures. Then, the canonical (NVT) ensemble simulations with the Nosé-Hoover thermostat were utilized to ascertain the lithium diffusion properties within c-LLZO. A time step of 2 fs was used for all NpT and NVT simulations. In the case of NVT simulations, following a 100 ps equilibration period, 2-ns trajectories’ mean squared displacements of Li ions were employed to calculate the self-diffusion coefficients at various target temperatures. Additionally, we conducted an uncertainty analysis of the diffusion coefficients, applying the empirical error estimation method as detailed in reference41.
To construct the phase diagram of the $\rm BaTiO_{3}$ crystal structure, a supercell consisting of a 10×10×10 $\rm R3m$ $\rm BaTiO_{3}$ lattice with 5000 atoms was created. The crystal structure data for $\rm BaTiO_{3}$ was sourced from the Materials Project36. NpT simulations with the Nosé-Hoover thermostat and Parrinello-Rahman barostat were performed with a 1-fs timestep. Throughout the simulations, the pressure was consistently maintained at 1 atmosphere. To regulate temperature, the procedure began at 5 K and involved incrementing the temperature by 10 K every 20 ps. For the comprehensive calculation method of the local polarization, please refer to the Supplementary Information.
@@
We utilized a 13760-atom structure consisting of a ca. 13.6-nm thick $\rm Im\bar{3}m$ lithium metal anode with [100] facet, paired with a ca. 19.1 nm thick $\rm Pnma$ $\beta-\rm{Li_{3}PS_{4}}$ electrolyte with [010] facet to illustrate the SEI formation across realistic spatial and temporal scales. The system was structured with periodic lateral dimensions measuring 3.1 nm by 2.6 nm. The unit cell of anode and electrolyte structures originated from the Materials Project36. The entire cell was subject to periodic boundary conditions, with a 1 nm atomic layer fixed on each side along the x-direction to ensure a singular interface reaction. NpT MD simulations were conducted at a standard temperature of 300 K and a pressure of 1 atm by Nosé-Hoover thermostat and Parrinello-Rahman barostat with a timestep of 2 fs. Initial velocities for the particles were assigned randomly following the Maxwell-Boltzmann distribution, and the simulations were carried out for a duration of 4 ns.
Data and software availability
@@
The dataset is accessible through the Materials Project36, 16. Detailed PQEq parameters are provided in the Supplementary Information. The JAX implementation code is available throught https://github.com/reaxnet/reaxnet.